1. Suppose the pmf of a random variable x is given by f(x)={1/6,for x=1,2,3,4,5,6

{0,elsewhere

Obtain the pmf of y=2x2 and z=x-3

2.Let the pmf of a random variable X be given by f(x)={(x2+1)/18 for x=0,1,2,3

{0,elsewhere

determine the pmf of Y=x2+1

3.Suppose the pmf of a random variable x is given by f(x)={x/10 for x=0,1,2,3,4

{0,elsewhere

Obtain the pmf of y=|x-2|

4.Let the pmf of a random variable x be given by f(x)={(1/2)x for x=1,2,3

{0,elsewhere

determine the pmf

y={1 if x is odd

{0 if x is even

5.Suppose the random variable x has the pmf is given by f(x)={1/6(5/6)x for x=0,1,2,3,....

{0,elsewhere

Obtain the pmf of y=x-1


1
Expert's answer
2021-07-12T13:26:55-0400

1.


"f(y) = \begin{cases}
 1/6 &\text{for } y=2,8,18,32,50,72 \\
 0 &\text{elsewhere } 
\end{cases}"


"f(z) = \begin{cases}
 1/6 &\text{for } z=-2,-1,0,1,2,3 \\
 0 &\text{elsewhere } 
\end{cases}"


2.


"f(y) = \begin{cases}
 y/18 &\text{for } y=1,2,5,10 \\
 0 &\text{elsewhere } 
\end{cases}"


3.


"\begin{matrix}
 x & 0 & 1 & 2 & 3 & 4 \\
 y & 2 & 1 & 0 & 2 & 2 
\end{matrix}"


"\begin{matrix}
 y & 0 & 1 & 2 \\
 p(y) & 2/10 & 4/10 & 4/10 
\end{matrix}"


"f(y) = \begin{cases}
 1/5 &\text{for } y=0 \\
 2/5 &\text{for } y=1 \\
 2/5 &\text{for } y=2 \\
 0 &\text{elsewhere } 
\end{cases}"

4.


"f(x) = \begin{cases}
 1/2^x &\text{for } x=1,2,3,4, ... \\
 0 &\text{elsewhere } 
\end{cases}"

"\dfrac{1}{2}+\dfrac{1}{8}+\dfrac{1}{32}+...=\displaystyle\sum_{i=1}^{\infin}\dfrac{1}{2^{n+1}}=\dfrac{1}{2}"


"f(y) = \begin{cases}
 1/2 &\text{for } y=0 \\
 1/2 &\text{for } y=1 \\
 0 &\text{elsewhere } 
\end{cases}"

5.


"f(x) = \begin{cases}
 \dfrac{5}{36}\bigg(\dfrac{5}{6}\bigg)^y &\text{for } y=-1,0,1,2,3,4, ... \\
 0 &\text{elsewhere } 
\end{cases}"